OpenStudy (anonymous):

Need help showing check. I've solved these two problems, if you can show me the work for how to check please. 2) 2x-3y=16 -3x+2y=-19 x=5, y=-2, Show check please

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Show how to check it? okay...well you got the answers that x = 5 and y = -2 so to check...all you do is plug those back into the original equations. 2x - 3y = 16 -3x + 2y = -19 Replace all the 'x' you see with 5 and then all the 'y' you see with -2 2(5) - 3(-2) = 16 10 - (-6) = 16 10 + 6 = 16 16 = 16 <--- so this is correct -3(5) + 2(-2) = -19 -15 + (-4) = -19 -15 - 4 = -19 -19 = -19 <--- this is also correct...and that is your check

OpenStudy (johnweldon1993):

Hope that helps :)

OpenStudy (anonymous):

-3x+y=-2 -2x+3y=15 x=3, y=7 Can you show me the work for how to solve for y? This is what I did for X -3x+y=-2 y=3x-2 -2x+3y=15 -2x+3(3x-2)=15 7x=21 x=3

OpenStudy (johnweldon1993):

Alright...to solve for 'y' you can either just plug in x = 3 into one of the original equations...or you can do that whole process again for 'y' if you want the latter...it would be -3x+y=-2 -2x+3y=15 Lets solve the top on for 'x' -3x + y = -2 -3x = -2 - y x = 2/3 + y/3 Now plug that into the second equation for 'x' -2x + 3y = 15 -2(2/3 + y/3) + 3y = 15 -4/3 - 2y/3 + 3y = 15 Lets multiply everything by 3 to make this easier -4 - 2y + 9y = 45 -4 + 7y = 45 Add 4 to both sides 7y = 49 Divide both sides by 7 y = 7

OpenStudy (anonymous):

Yes it does, thank you so much for your help!

OpenStudy (johnweldon1993):

Anytime! :)